OpenStudy (anonymous):

An arc of length 10 meters is formed by a central angle A on a circle of radius 4. The measure of A in degrees (to two decimal places) is _____. A. 216.76 B. 143.24 C. 2.50 D. 36.76

ganeshie8 (ganeshie8):

you may use these :- 1. \(\theta = \frac{l}{r}\) 2. \(\pi\ radians = 180\ degrees\)

OpenStudy (anonymous):

look at pic

OpenStudy (anonymous):

so B?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Thanks :)